On January 1, 2021, Essence Communications issued $600,000 of its 10-year, 10% bonds for $531,180. The bonds were priced to yield 12%. Interest is payable semiannually on June 30 and December 31. Essence Communications records interest at the effective rate and elected the option to report these bonds at their fair value. On December 31, 2021, the market interest rate for bonds of similar risk and maturity was 11%. The bonds are not traded on an active exchange. The decrease in the market interest rate was due to a 1% decrease in general (risk-free) interest rates. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required: 1. Using the information provided, estimate the fair value of the bonds at December 31, 2021. 2. to 4. Prepare the journal entries to record interest on June 30, 2021 (the first interest payment), on December 31, 2021 (the second interest payment) and to adjust the bonds to their fair value for presentation in the December 31, 2021, balance sheet.
Answer #1

Interest would be halved and number of period will be doubles because of semi annual interest payments

Please find below table useful to compute desired results: -

A B Par Value of Bonds Interest Rate Issued Price Yield Market Interest Rate 600000 =10/100 531180 =12/100 =11/100 Interest 1

End results would be as follows: -

D B Par Value of Bonds Interest Rate Issued Price Yield Market Interest Rate 600,000 10% 531,180 12% 11% Interest 1. Fair Val